OpenStudy (anonymous):

A simple harmonic oscillator in the physics lab is measured to have an amplitude of 180.00 cm and a maximum speed of 350.0 cm/s. What is its speed when the displacement is 3.40 cm?

OpenStudy (lastdaywork):

Amplitude = A Max Speed = Aω Displacement = x Use Equations x=A sin(ωt) v=Aω cos(ωt)

OpenStudy (anonymous):

i tried and couldnt solve it

OpenStudy (lastdaywork):

Where did you got struck?

OpenStudy (lastdaywork):

If you are struggling with trigonometry, use - \[\sin \theta=\sqrt{1-\cos^{2} \theta}\] It is not true in general, but will work in this case.

OpenStudy (loser66):

we have maxspeed = A \(\omega\) ---> \(\omega = \dfrac{maxspeed}{A}\)= \(\dfrac{350}{180}=1.94\) And we have formula to calculate velocity at any time: \[v = \sqrt{\omega(A^2 -x^2)}\] just replace number in.

OpenStudy (anonymous):

I didnt get the right answer with that either